Rank the following mutually exclusive projects based on profitability index (PI). Cost of capital = 6%.

Project Initial Investment Cash flow 1 Cash flow 2
A 500 300 250
B 600 450 300
C 450 300 300
Answer #1

Project A:

Present Value of Cash Inflows = 300/1.06 + 250/1.06^2
Present Value of Cash Inflows = 505.5180

Profitability Index = Present Value of Cash Inflows / Initial Investment
Profitability Index = 505.5180 / 500
Profitability Index = 1.01

Project B:

Present Value of Cash Inflows = 450/1.06 + 300/1.06^2
Present Value of Cash Inflows = 691.5272

Profitability Index = Present Value of Cash Inflows / Initial Investment
Profitability Index = 691.5272 / 600
Profitability Index = 1.15

Project C:

Present Value of Cash Inflows = 300/1.06 + 300/1.06^2
Present Value of Cash Inflows = 550.0178

Profitability Index = Present Value of Cash Inflows / Initial Investment
Profitability Index = 550.0178 / 450
Profitability Index = 1.22

Rank:
I = Project C
II = Project B
III = Project A